### v3

@@ 2.4 Phase 3 - Conducting a Survey
A questionnaire-based survey approach is used to evaluate the taxonomies of issues, causes, and solutions based on mining developer discussions and conducting practitioner interviews. We adopted Kitchenham and Pfleeger’s guidelines for conducting surveys [29] and used an anonymous survey to increase response rates [30].
@@ 2.4.2 Conducting the Web-based Survey
We adopted a cross-sectional survey design, which is appropriate for collecting information at one given point in time across a sample population [29]. Surveys can be conducted in many ways, such as Web-based online questionnaires and phone surveys [31]. We decided to conduct a Web-based survey because these surveys can help to (i) minimize the time and cost, (ii) collect the responses from geographically distributed respondents, (iii) minimize time zone constraints, and (iv) save the effort of researchers to collect data in a textual, graphical, or structured format [31]. To document different types of responses while maintaining the granularity of information, we structured the questionnaire into a total of 12 questions organized under four sections (see the Survey Questionnaire sheet in [19]).
@@
Microservices Practitioners’ Perspective: To evaluate the taxonomies of issues, causes, and solutions, we asked six survey questions (both Likert scale and open-ended, see the Survey Questionnaire sheet in [19]) from microservices practitioners. We provided a list of 19 issue categories and asked survey participants to respond to each category on a 5-point Likert scale (Very Often, Often, Sometimes, Rarely, Never). Similarly, regarding causes, we provided 8 categories and asked practitioners to respond to each category on a 5-point Likert scale (Strongly Agree, Agree, Neutral, Disagree, Strongly Disagree). We also provided 8 categories of solutions and asked practitioners to respond to each category on a 5-point Likert scale. Along with these 5-point Likert scales, we added one option to know the familiarity of the survey respondents with the listed categories. We also asked three open-ended questions to identify the missing issue, causes, and solutions in the provided categories. All the open-ended responses were further analyzed through thematic classification and adjusted in the taxonomies.
@@ 2.4.3 Data Analysis
[…]
@@ 3 Results
This section presents the analyzed results of this study, addressing the four RQs outlined in Section 2.1. The analyzed results are further organized as categories (e.g., Technical Debt), subcategories (e.g., Code Debt), and types (e.g., code smell). We present categories in boldface, subcategories in italic, and types in small capitals. The relevant examples are provided as quoted messages along with their issue ID numbers to facilitate the traceability to our dataset (see the Initial Codes sheet in [19]). We report the types of issues in Section 3.1 (see Figure 4), the types of causes in Section 3.2 (see Table 3), the types of solutions in Section 3.3 (see Table 4), and practitioners’ perspective on these taxonomies in Section 3.4 (see Table 5).
@@ 3.1 Types of Issues (RQ1)
1. Technical Debt (687/2698, 25.46%): Technical Debt (TD) is “a metaphor reflecting technical compromises that can yield short-term benefit but may hurt the long-term health of a software system” [33]. This is the largest category in the taxonomy of microservices issues and includes a wide range of issues related to Code Debt and Service Design Debt. The interviewees also mentioned several issues regarding this category, and one representative quotation is depicted below.
[…]
@@
We identified and classified 19 types of TD issues in 2 subcategories (see Figure 4 and the Issue Taxonomy sheet in [19]). Each of them is briefly described below.
@@
We identified and classified 55 types of CI/CD issues in 7 subcategories (see Figure 4 and the Issue Taxonomy sheet in [19]). Each of them is briefly described below.
@@
3. Exception Handling Issue (228/2698, 8.45%): Exception handling is used to respond the unexpected errors during the running state of software systems (e.g., microservices systems), and it helps to avoid the software system being crashed unexpectedly. This category represents the issues practitioners face when handling various kinds of exceptions in microservices systems. We identified and classified 44 types of Exception Handling issues in 5 subcategories (see Figure 4 and the Issue Taxonomy sheet in [19]). Each of them is briefly described below.
@@
We identified and classified 34 types of service execution and communication issues in 3 subcategories (see Figure 4 and the Issue Taxonomy sheet in [19]). Each of them is briefly described below
@@
We identified and classified 37 types of security issues in 4 subcategories (see Figure 4 and the Issue Taxonomy sheet in [19]). Each of them is briefly described below.
@@
6. Build Issue (210/2698, 7.78%): Build is a process of preparing an application program for software release by collecting and compiling all required source files. The outcome of this process could be several types of artifacts, such as binaries and executable programs. We identified and classified 20 types of build issues in 3 subcategories (see Figure 4 and the Issue Taxonomy sheet in [19]). Each of them is briefly described below.
@@
We identified and classified 16 types of configuration issues in 2 subcategories (see Figure 4 and the Issue Taxonomy sheet in [19]). Each of them is briefly reported below.
@@
We identified and classified 17 types of monitoring issues in 3 subcategories (see Figure 4 and the Issue Taxonomy sheet in [19]). Each of them is briefly described below.
@@
9. Compilation Issue (79/2698, 2.92%): This category reports compilation errors, which mainly occur when the compiler cannot compile source code due to errors in the code or errors with the compiler itself. We identified and classified 9 types of compilation issues in 2 subcategories (see Figure 4 and the Issue Taxonomy sheet in [19]). Each of them is briefly described below.
@@
We identified and classified 20 types of testing issues in 3 subcategories (see Figure 4 and the Issue Taxonomy sheet in [19]). Each of them is briefly described below.
@@
11. Documentation Issue (75/2698, 2.77%): Documentation for microservices systems may suffer from several problems. We identified and classified 8 types of documentation issues in two subcategories (see Figure 4 and the Issue Taxonomy sheet in [19]). Each of them is briefly described below.
@@
12. Graphical User Interface (GUI) Issue (70/2698, 2.59%): This category reports the problems that can wreck the GUI of microservices systems. We identified and classified 68 types of GUI issues in 3 subcategories (see Figure 4 and the Issue Taxonomy sheet in [19]). Each of them is briefly described below.
@@
13. Update and Installation Issue (68/2698, 2.52%): This category gathers the identified problems related to update and installation of the packages, libraries, tools, and containerization platforms required to develop and manage microservices systems. We identified and classified 16 types of update and installation issues in 2 subcategories (see Figure 4 and the Issue Taxonomy sheet in [19]). Each of them is briefly described below.
@@
We identified and classified 24 types of database issues in 3 subcategories (see Figure 4 and the Issue Taxonomy sheet in [19]). Each of them is briefly described below
@@
15. Storage Issue (54/2698, 2.00%): This category reports storage space problems during the development, execution, and management of microservices systems. We identified and classified 13 types of storage issues in 2 subcategories (see Figure 4 and the Issue Taxonomy sheet in [19]), and each subcategory is further described below.
@@
We identified and classified 16 types of performance issues in three subcategories (see Figure 4 and the Issue Taxonomy sheet in [19]). Each of them is briefly described below.
@@
17. Networking Issue (41/2698, 1.51%): Deploying, executing, and communicating in microservices systems over the network is complex. It is observed that many problems may disrupt the network. We identified and classified 20 types of networking issues in 2 subcategories (see Figure 4 and the Issue Taxonomy sheet in [19]). Each of them is briefly described below.
@@
18. Typecasting Issue (35/2698, 1.29%): This category is related to the typecasting issues that occur when assigning a value of one primitive data type to another type. We identified and classified 9 types of typecasting issues in 2 subcategories (see Figure 4 and the Issue Taxonomy sheet in [19]). Each of them is briefly described below.
@@
19. Organizational Issue (7/2698, 0.25%): We derived this category based on the interviewees’ feedback on the taxonomy of the issues (see Figure 4 and the Issue Taxonomy sheet in [19]). The interviewees only mentioned three types of issues in this category, including team management, operational and tooling overhead, and service size. We depicted two representative quotations below.
[…]
@@ 3.2 Causes of Issues (RQ2)
The taxonomy of causes of microservices issues is provided in Table 3. It is worth mentioning that not all the issue discussions provide the information about their causes. Therefore, we identified 2,225 issue discussions containing information about the causes. The taxonomy of causes is derived by mining developer discussions (i.e., 2,225 instances of causes), conducting practitioner interviews (i.e., 31 instances of causes, see Section 2.3.3), and conducting a survey (i.e., 11 instances of causes, see Section 3.4). Hence, we got a total of 2,267 instances of causes. We identified a total of 228 types of causes that can be classified into 8 categories and 26 subcategories. Due to space limitations, we only list the top two types of causes for each subcategory in Table 3. The detail of the types of causes can be found in the dataset [19]. The results show that General Programming Error (860 out of 2267), Missing Features and Artifacts (386 out of 2267), and Invalid Configuration & Communication Problems (382 out of 2267) are the top three categories of causes. Each cause category is briefly discussed below.
1. General Programming Error (GPE) (860/2267, 37.93%): This category captures the causes that are based on a broad range of errors occurred in different phases of microservices system development, such as coding (e.g., syntax errors), testing (e.g., incorrect test cases), and maintenance (e.g., wrong examples in documentation) phases. We identified and classified 78 types of GPE causes in six subcategories (see Table 3 and the Cause Taxonomy sheet in [19]). Each of them is briefly described below.
@@
2. Missing Features and Artifacts (MFA) (386/2267, 17.02%): This category represents the causes behind microservices systems issues that occur due to missing required features, packages, files, variables, and documentation and tool support. We identified and classified 27 types of MFA causes in four subcategories (see Table 3 and the Cause Taxonomy sheet in [19]). Each of them is briefly described below.
@@
We identified and classified 29 types of causes in two subcategories (see Table 3 and the Cause Taxonomy sheet in [19]). Each of them is briefly described below.
@@
We identified and classified 28 types of causes in five subcategories (see Table 3 and the Cause Taxonomy sheet in [19]). Each of them is briefly described below.
  • Compatibility and Dependency (59, 2.60%): A typical microservices system consists of several independent services running on multiple servers or hosts [34]. However, some microservices also depend on other microservices to complete business operations. Usually, practitioners ensure the compatibility (e.g., backward compatibility) of each microservice with previous versions of the microservices systems during the upgrading. In this subcategory, we identified 5 types of causes, in which the leading three types of causes are compatibility error, and outdated dependency.
@@
We identified and classified 17 types of SD&IA causes in 2 subcategories (see Table 3 and the Cause Taxonomy sheet in [19]). Each of them is briefly described below.
@@
We identified and classified 21 types of PSM causes in 3 subcategories (see Table 3 and the Cause Taxonomy sheet in [19]). Each of them is briefly described below.
@@
7. Insufficient Resources (IR) (93/2267, 4.10% ): Microservices systems are at risk of delivering the required outcome without sufficient resources. We identified and classified 4 types of IR causes in two subcategories (see Table 3 and the Cause Taxonomy sheet in [19]). Each of them is briefly described below.
@@
8. Fragile Code (FC) (24/2267, 1.05%): Fragile code refers to code that is difficult to change, and a minor modification in fragile code may break the service or module. We identified and classified 6 types of FC causes in two subcategories (see Table 3 and the Cause Taxonomy sheet in [19]). Each of them is briefly described below.

### v4

@@ 2.4 Phase 3 - Conducting a Survey
A questionnaire-based survey approach is used to evaluate the taxonomies of issues, causes, and solutions based on mining developer discussions and conducting practitioner interviews. We adopted Kitchenham and Pfleeger’s guidelines for conducting surveys [kitchenham2008personal] and used an anonymous survey to increase response rates [tyagi1989effects].
@@ 2.4.2 Conducting the Web-based Survey
We adopted a cross-sectional survey design, which is appropriate for collecting information at one given point in time across a sample population [kitchenham2008personal]. Surveys can be conducted in many ways, such as Web-based online questionnaires and phone surveys [lethbridge2005studying]. We decided to conduct a Web-based survey because these surveys can help to (i) minimize the time and cost, (ii) collect the responses from geographically distributed respondents, (iii) minimize time zone constraints, and (iv) save the effort of researchers to collect data in a textual, graphical, or structured format [lethbridge2005studying]. To document different types of responses while maintaining the granularity of information, we structured the questionnaire into a total of 12 questions organized under four sections (see the Survey Questionnaire sheet in [replpack]).
@@
Microservices Practitioners’ Perspective: To evaluate the taxonomies of issues, causes, and solutions, we asked six survey questions (both Likert scale and open-ended, see the Survey Questionnaire sheet in [replpack]) from microservices practitioners. We provided a list of 19 issue categories and asked survey participants to respond to each category on a 5-point Likert scale (Very Often, Often, Sometimes, Rarely, Never). Similarly, regarding causes, we provided 8 categories and asked practitioners to respond to each category on a 5-point Likert scale (Strongly Agree, Agree, Neutral, Disagree, Strongly Disagree). We also provided 8 categories of solutions and asked practitioners to respond to each category on a 5-point Likert scale. Along with these 5-point Likert scales, we added one option to know the familiarity of the survey respondents with the listed categories. We also asked three open-ended questions to identify the missing issue, causes, and solutions in the provided categories. All the open-ended responses were further analyzed through thematic classification and adjusted in the taxonomies.
@@ 2.4.3 Data Analysis
We used descriptive statistics and constant comparison techniques [hoda2017becoming] to analyze the quantitative (i.e., closed-ended questions) and qualitative (i.e., open-ended questions) responses to the survey questions, respectively. To better understand practitioners’ perspectives through Likert answers on issues, causes, and solutions in microservices systems, we used the Wilcoxon rank-sum test, i.e., participants who have $\leq$ 6 years of experience (49 responses) versus participants who have $\geq$ 6 years of experience with microservices systems (101 responses). We used 6 years as a breaking point for separating the groups because 6 years is almost in the middle of practitioners’ experience with microservices systems. We used the  symbol to indicate a significant difference between participant groups who have Experience $\leq$ 6 years vs. Experience $\geq$ 6 years. To analyze the results of open-ended questions, we employed the constant comparison method from grounded theory [hoda2017becoming], which allowed us to systematically identify and understand various issues, causes, and solutions from practitioners’ perspective. We began the process by coding the responses and assigning relevant labels. Throughout the analysis, we persistently compared new responses with existing codes (i.e., types of issues, causes, and solutions from the taxonomies) and made refinements as needed.
@@ 3 Results
This section presents the analyzed results of this study, addressing the four RQs outlined in Section 2.1. The analyzed results are further organized as categories (e.g., Technical Debt), subcategories (e.g., Code Debt), and types (e.g., code smell). We present categories in boldface, subcategories in italic, and types in small capitals. The relevant examples are provided as quoted messages along with their issue ID numbers to facilitate the traceability to our dataset (see the Initial Codes sheet in [replpack]). We report the types of issues in Section 3.1 (see Figure 4), the types of causes in Section 3.2 (see Table 3), the types of solutions in Section 3.3 (see Table 4), and practitioners’ perspective on these taxonomies in Section 3.4 (see Table 5).
@@ 3.1 Types of Issues (RQ1)
1. Technical Debt (687/2698, 25.46%): Technical Debt (TD) is “a metaphor reflecting technical compromises that can yield short-term benefit but may hurt the long-term health of a software system” [li2015systematic]. This is the largest category in the taxonomy of microservices issues and includes a wide range of issues related to Code Debt and Service Design Debt. The interviewees also mentioned several issues regarding this category, and one representative quotation is depicted below.
@@
We identified and classified 19 types of TD issues in 2 subcategories (see Figure 4 and the Issue Taxonomy sheet in [replpack]). Each of them is briefly described below.
@@
We identified and classified 55 types of CI/CD issues in 7 subcategories (see Figure 4 and the Issue Taxonomy sheet in [replpack]). Each of them is briefly described below.
@@
3. Exception Handling Issue (228/2698, 8.45%): Exception handling is used to respond the unexpected errors during the running state of software systems (e.g., microservices systems), and it helps to avoid the software system being crashed unexpectedly. This category represents the issues practitioners face when handling various kinds of exceptions in microservices systems. We identified and classified 44 types of Exception Handling issues in 5 subcategories (see Figure 4 and the Issue Taxonomy sheet in [replpack]). Each of them is briefly described below.
@@
We identified and classified 34 types of service execution and communication issues in 3 subcategories (see Figure 4 and the Issue Taxonomy sheet in [replpack]). Each of them is briefly described below
@@
We identified and classified 37 types of security issues in 4 subcategories (see Figure 4 and the Issue Taxonomy sheet in [replpack]). Each of them is briefly described below.
@@
6. Build Issue (210/2698, 7.78%): Build is a process of preparing an application program for software release by collecting and compiling all required source files. The outcome of this process could be several types of artifacts, such as binaries and executable programs. We identified and classified 20 types of build issues in 3 subcategories (see Figure 4 and the Issue Taxonomy sheet in [replpack]). Each of them is briefly described below.
@@
We identified and classified 16 types of configuration issues in 2 subcategories (see Figure 4 and the Issue Taxonomy sheet in [replpack]). Each of them is briefly reported below.
@@
We identified and classified 17 types of monitoring issues in 3 subcategories (see Figure 4 and the Issue Taxonomy sheet in [replpack]). Each of them is briefly described below.
@@
9. Compilation Issue (79/2698, 2.92%): This category reports compilation errors, which mainly occur when the compiler cannot compile source code due to errors in the code or errors with the compiler itself. We identified and classified 9 types of compilation issues in 2 subcategories (see Figure 4 and the Issue Taxonomy sheet in [replpack]). Each of them is briefly described below.
@@
We identified and classified 20 types of testing issues in 3 subcategories (see Figure 4 and the Issue Taxonomy sheet in [replpack]). Each of them is briefly described below.
@@
11. Documentation Issue (75/2698, 2.77%): Documentation for microservices systems may suffer from several problems. We identified and classified 8 types of documentation issues in two subcategories (see Figure 4 and the Issue Taxonomy sheet in [replpack]). Each of them is briefly described below.
@@
12. Graphical User Interface (GUI) Issue (70/2698, 2.59%): This category reports the problems that can wreck the GUI of microservices systems. We identified and classified 68 types of GUI issues in 3 subcategories (see Figure 4 and the Issue Taxonomy sheet in [replpack]). Each of them is briefly described below.
@@
13. Update and Installation Issue (68/2698, 2.52%): This category gathers the identified problems related to update and installation of the packages, libraries, tools, and containerization platforms required to develop and manage microservices systems. We identified and classified 16 types of update and installation issues in 2 subcategories (see Figure 4 and the Issue Taxonomy sheet in [replpack]). Each of them is briefly described below.
@@
We identified and classified 24 types of database issues in 3 subcategories (see Figure 4 and the Issue Taxonomy sheet in [replpack]). Each of them is briefly described below
@@
15. Storage Issue (54/2698, 2.00%): This category reports storage space problems during the development, execution, and management of microservices systems. We identified and classified 13 types of storage issues in 2 subcategories (see Figure 4 and the Issue Taxonomy sheet in [replpack]), and each subcategory is further described below.
@@
We identified and classified 16 types of performance issues in three subcategories (see Figure 4 and the Issue Taxonomy sheet in [replpack]). Each of them is briefly described below.
@@
17. Networking Issue (41/2698, 1.51%): Deploying, executing, and communicating in microservices systems over the network is complex. It is observed that many problems may disrupt the network. We identified and classified 20 types of networking issues in 2 subcategories (see Figure 4 and the Issue Taxonomy sheet in [replpack]). Each of them is briefly described below.
@@
18. Typecasting Issue (35/2698, 1.29%): This category is related to the typecasting issues that occur when assigning a value of one primitive data type to another type. We identified and classified 9 types of typecasting issues in 2 subcategories (see Figure 4 and the Issue Taxonomy sheet in [replpack]). Each of them is briefly described below.
@@
19. Organizational Issue (7/2698, 0.25%): We derived this category based on the interviewees’ feedback on the taxonomy of the issues (see Figure 4 and the Issue Taxonomy sheet in [replpack]). The interviewees only mentioned three types of issues in this category, including team management, operational and tooling overhead, and service size. We depicted two representative quotations below.
@@ 3.2 Causes of Issues (RQ2)
The taxonomy of causes of microservices issues is provided in Table 3. It is worth mentioning that not all the issue discussions provide the information about their causes. Therefore, we identified 2,225 issue discussions containing information about the causes. The taxonomy of causes is derived by mining developer discussions (i.e., 2,225 instances of causes), conducting practitioner interviews (i.e., 31 instances of causes, see Section 2.3.3), and conducting a survey (i.e., 11 instances of causes, see Section 3.4). Hence, we got a total of 2,267 instances of causes. We identified a total of 228 types of causes that can be classified into 8 categories and 26 subcategories. Due to space limitations, we only list the top two types of causes for each subcategory in Table 3. The detail of the types of causes can be found in the dataset [replpack]. The results show that General Programming Error (860 out of 2267), Missing Features and Artifacts (386 out of 2267), and Invalid Configuration & Communication Problems (382 out of 2267) are the top three categories of causes. Each cause category is briefly discussed below.
1. General Programming Error (GPE) (860/2267, 37.93%): This category captures the causes that are based on a broad range of errors occurred in different phases of microservices system development, such as coding (e.g., syntax errors), testing (e.g., incorrect test cases), and maintenance (e.g., wrong examples in documentation) phases. We identified and classified 78 types of GPE causes in six subcategories (see Table 3 and the Cause Taxonomy sheet in [replpack]). Each of them is briefly described below.
@@
2. Missing Features and Artifacts (MFA) (386/2267, 17.02%): This category represents the causes behind microservices systems issues that occur due to missing required features, packages, files, variables, and documentation and tool support. We identified and classified 27 types of MFA causes in four subcategories (see Table 3 and the Cause Taxonomy sheet in [replpack]). Each of them is briefly described below.
@@
We identified and classified 29 types of causes in two subcategories (see Table 3 and the Cause Taxonomy sheet in [replpack]). Each of them is briefly described below.
@@
We identified and classified 28 types of causes in five subcategories (see Table 3 and the Cause Taxonomy sheet in [replpack]). Each of them is briefly described below.
  • Compatibility and Dependency (59, 2.60%): A typical microservices system consists of several independent services running on multiple servers or hosts [waseem2022]. However, some microservices also depend on other microservices to complete business operations. Usually, practitioners ensure the compatibility (e.g., backward compatibility) of each microservice with previous versions of the microservices systems during the upgrading. In this subcategory, we identified 5 types of causes, in which the leading three types of causes are compatibility error, and outdated dependency.
@@
We identified and classified 17 types of SD&IA causes in 2 subcategories (see Table 3 and the Cause Taxonomy sheet in [replpack]). Each of them is briefly described below.
@@
We identified and classified 21 types of PSM causes in 3 subcategories (see Table 3 and the Cause Taxonomy sheet in [replpack]). Each of them is briefly described below.
@@
7. Insufficient Resources (IR) (93/2267, 4.10% ): Microservices systems are at risk of delivering the required outcome without sufficient resources. We identified and classified 4 types of IR causes in two subcategories (see Table 3 and the Cause Taxonomy sheet in [replpack]). Each of them is briefly described below.
@@
8. Fragile Code (FC) (24/2267, 1.05%): Fragile code refers to code that is difficult to change, and a minor modification in fragile code may break the service or module. We identified and classified 6 types of FC causes in two subcategories (see Table 3 and the Cause Taxonomy sheet in [replpack]). Each of them is briefly described below.